# Finite–slope universal Rankin–Selberg $p$-adic $L$-function

### v1

@@ 2.3. Finiteness and base change for $(\varphi,\Gamma)$–cohomology
    The Iwasawa cohomology complex $C^{\bullet}_{\psi}(\mathsf{M})$ lies in $D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}^{\infty}_{A}(\Gamma)\bigr)$, and the $(\varphi,\Gamma)$–cohomology complex $C^{\bullet}_{\varphi,\Gamma}(\mathsf{M})$ lies in $D^{-}_{\mathrm{perf}}(A)$ [19, Thm. 4.4.1, Thm. 4.4.2].
@@
    $C^{\bullet}_{\psi}(\mathsf{M})\;\otimes^{\mathbf{L}}_{\mathcal{R}^{\infty}_{A}(\Gamma)}\mathcal{R}^{\infty}_{B}(\Gamma)\longrightarrow C^{\bullet}_{\psi}\bigl(\mathsf{M}\widehat{\otimes}_{A}B\bigr),$
@@
We apply Theorem 5 with $A=\mathscr{O}(U)$ and $\mathsf{M}=\mathscr{D}$ or $\mathsf{M}=\mathscr{D}^{+}$. This shows that
$C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}),\;C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}^{+})\;\in\;D^{-}_{\mathrm{perf}}\bigl(\mathscr{O}(U)\bigr),$
@@
$C^{\bullet}_{\psi}(\mathscr{D}),\;C^{\bullet}_{\psi}(\mathscr{D}^{+})\;\in\;D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}^{\infty}_{\mathscr{O}(U)}(\Gamma)\bigr).$
Moreover, for any morphism of affinoid algebras $\mathscr{O}(U)\to B$ (for instance $B=\mathscr{O}(U^{\prime})$ for an affinoid subdomain $U^{\prime}\subset U$, or $B=k(x)$ for the residue field at a point $x\in U$), the base–change isomorphisms of Theorem 5(ii) give quasi-isomorphisms
$C^{\bullet}_{\varphi,\Gamma}(\mathscr{D})\otimes^{\mathbf{L}}_{\mathscr{O}(U)}B\;\xrightarrow{\ \sim\ }\;C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}\widehat{\otimes}_{\mathscr{O}(U)}B),$
@@
$C^{\bullet}_{\psi}(\mathscr{D})\otimes^{\mathbf{L}}_{\mathcal{R}^{\infty}_{\mathscr{O}(U)}(\Gamma)}\mathcal{R}^{\infty}_{B}(\Gamma)\;\xrightarrow{\ \sim\ }\;C^{\bullet}_{\psi}\bigl(\mathscr{D}\widehat{\otimes}_{\mathscr{O}(U)}B\bigr),$
@@
Remark 3.
For each $x\in U$, the fibre of $C^{\bullet}_{\varphi,\Gamma}(\mathscr{D})$ at $x$ computes the $(\varphi,\Gamma)$–cohomology of $D_{\operatorname{rig}}^{\dagger}(V_{x})$, and similarly for $\mathscr{D}^{+}$. Likewise, the fibres of $C^{\bullet}_{\psi}(\mathscr{D})$ and $C^{\bullet}_{\psi}(\mathscr{D}^{+})$ compute the Iwasawa cohomology of $D_{\operatorname{rig}}^{\dagger}(V_{x})$ and $D_{\operatorname{rig}}^{\dagger}(V_{x}^{+})$, respectively. Thus these complexes give a uniform description, over $U$, of the local Galois cohomology groups and the Bloch–Kato local conditions at $p$ for the specialisations $V_{x}$ and $V_{x}^{+}$ in the sense of [16, §3]. This is the cohomological input needed in Step 3 to construct a big logarithm/regulator map
$\mathcal{L}_{V,V^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V^{\ast}(1))\;\longrightarrow\;\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Z}_{p}}\mathscr{O}(U),$
and to relate its specialisations at classical points to the Bloch–Kato exponentials appearing in the Bloch–Kato Selmer formalism of [16, §3].
Step 3: Construction of the big logarithm and of the finite–slope Rankin–Selberg regulator
In this step we fix the affinoid neighbourhood $U$ and the Panchishkin family
$(V_{U},V_{U}^{+})$
from Steps 1 and 2. Thus $V_{U}$ is a rank–$4$ finite projective $\mathcal{O}(U)$-module equipped with a continuous $\mathcal{O}(U)$-linear action of $G_{\mathbb{Q}}$, and for each point $x\in U$ the specialisation $V_{x}$ is de Rham at $p$ and equipped with a Panchishkin subspace $V_{x}^{+}\subset V_{x}$ of fixed rank $r$.
We first introduce Iwasawa–theoretic notation and recall the local Perrin–Riou big logarithm for a single Panchishkin representation. We then explain how to extend it to our family $(V_{U},V_{U}^{+})$ using the relative $(\varphi,\Gamma)$–module technology of [19] and the Perrin–Riou formalism in the guise of [18, 24, 25, 26, 15]. Finally we apply this family big logarithm to the Beilinson–Flach Euler system of [21, 25, 15] to define the finite–slope universal Rankin–Selberg pp–adic LL–function $L_{p}^{\mathrm{fs}}$ and prove its interpolation property.
(3.1) Iwasawa cohomology and the distribution algebra.
@@
denote the algebra of $\mathbb{Q}_{p}$–valued locally analytic distributions on $\Gamma$, i.e. the completion of $\Lambda_{\mathbb{Q}_{p}}(\Gamma)$ in its natural Fréchet topology (cf. [29, §2.2] and [30, §2.1]). Equivalently, $\mathcal{H}(\Gamma)$ is the strong continuous dual of the space of locally analytic $\mathbb{Q}_{p}$–valued functions on $\Gamma$.
For any pp–adic representation $W$ of $G_{\mathbb{Q}_{p}}$ on a finite-dimensional $\mathbb{Q}_{p}$–vector space, its (cyclotomic) Iwasawa cohomology is defined by
$H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W):=H^{i}\bigl(\mathbb{Q}_{p},W\otimes_{\mathbb{Q}_{p}}\Lambda(\Gamma)^{\vee}\bigr)\qquad(i\geq 0),$
where $\Lambda(\Gamma)^{\vee}:=\operatorname{Hom}_{\mathrm{cts}}(\Lambda(\Gamma),\mathbb{Q}_{p})$ with the natural $G_{\mathbb{Q}_{p}}$–action; see [18, App. A.2–A.4] and Greenberg [28] for this definition and its basic properties. By construction, $\Lambda(\Gamma)$ acts on $\Lambda(\Gamma)^{\vee}$ via the right regular representation, and hence $H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)$ carries a natural structure of $\Lambda(\Gamma)$-module for every $i$.
Lemma 6 (Finiteness of local Iwasawa cohomology).
If $W$ is a finite-dimensional $\mathbb{Q}_{p}$–representation of $G_{\mathbb{Q}_{p}}$, then $H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)$ is a finitely generated $\Lambda(\Gamma)$–module for all $i\geq 0$. In particular, $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)$ is finitely generated over $\Lambda(\Gamma)$.
@@
Choose a $G_{\mathbb{Q}_{p}}$–stable $\mathbb{Z}_{p}$–lattice $T\subset W$, and write $K_{\infty}=\mathbb{Q}_{p}(\mu_{p^{\infty}})$. Following [18, App. A.2], let
@@
with respect to the corestriction maps; then $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ is naturally a $\Lambda(\Gamma)$–module, and there is a canonical identification
$Z^{i}_{\infty}(\mathbb{Q}_{p},T)\;\cong\;H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},T):=H^{i}\bigl(\mathbb{Q}_{p},T\otimes_{\mathbb{Z}_{p}}\Lambda(\Gamma)^{\vee}\bigr)$
for all $i$ (see [18, App. A.2–A.3]). Moreover, Proposition A.2.3 of loc. cit. shows that $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ is a finitely generated $\Lambda(\Gamma)$–module for $i=0,1,2$, and $Z^{i}_{\infty}(\mathbb{Q}_{p},T)=0$ for $i\geq 3$. Since $W=T\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$, we have
$H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)\;\cong\;H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},T)\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}\;\cong\;Z^{i}_{\infty}(\mathbb{Q}_{p},T)\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$
as $\Lambda(\Gamma)$–modules, and the finite generation of $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ implies the finite generation of $H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)$ for all $i$. ∎
In order to apply Perrin–Riou’s regulator, we pass from $\Lambda(\Gamma)$–coefficients to the algebra $\mathcal{H}(\Gamma)$ of locally analytic distributions by scalar extension along the canonical map $\Lambda(\Gamma)\to\mathcal{H}(\Gamma)$ and set
$H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)_{\mathcal{H}}\;:=\;H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)\,\widehat{\otimes}_{\Lambda(\Gamma)}\,\mathcal{H}(\Gamma).$
When no confusion can arise, we continue to denote this $\mathcal{H}(\Gamma)$-module simply by $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W)$.
If $A$ is a $\mathbb{Q}_{p}$–affinoid algebra and $W_{A}$ is a finite projective AA–module with a continuous AA–linear $G_{\mathbb{Q}_{p}}$–action, we define the family Iwasawa cohomology by
$H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{A}):=H^{i}\bigl(\mathbb{Q}_{p},W_{A}\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)^{\vee}\bigr)\qquad(i\geq 0);$
this is a priori an $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$–module. Let $\mathcal{R}_{A}^{\infty}(\Gamma)$ denote the Fréchet–Stein Iwasawa algebra of $\Gamma$ over $A$ considered in [19, §4.2]; there is a natural finite flat homomorphism
@@
The relative $(\varphi,\Gamma)$–module formalism of Kedlaya–Pottharst–Xiao [19] gives the following more precise statement.
Proposition 2 (Relative local Iwasawa cohomology via $(\varphi,\Gamma)$–modules).
Let $A$ be a reduced $\mathbb{Q}_{p}$–affinoid algebra and $W_{A}$ a finite projective AA–module with a continuous AA–linear $G_{\mathbb{Q}_{p}}$–action. Let $\mathscr{D}_{A}:=D_{\operatorname{rig}}^{\dagger}(W_{A})$ be the associated family of $(\varphi,\Gamma)$–modules over the relative Robba ring $\mathcal{R}_{A}$. Then:
@@
    and
    $H^{i}\bigl(C^{\bullet}_{\psi}(\mathscr{D}_{A})\bigr)\;\cong\;H^{i}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{A})\,\widehat{\otimes}_{\Lambda(\Gamma)}\mathcal{R}_{A}^{\infty}(\Gamma)$
@@
    in particular, $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{A})$ is a finite projective $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$–module, and for every morphism of affinoids $A\to B$ the canonical map
    $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{A})\,\widehat{\otimes}_{A\widehat{\otimes}\Lambda(\Gamma)}B\widehat{\otimes}\Lambda(\Gamma)\xrightarrow{\ \sim\ }H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{B})$
@@
The existence, perfectness, and base-change properties of the complexes $C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{A})$ and $C^{\bullet}_{\psi}(\mathscr{D}_{A})$, as well as the identifications of their cohomology with Galois and Iwasawa cohomology after extension of scalars to $\mathcal{R}_{A}^{\infty}(\Gamma)$, are proved in [19, Rem. 4.3.3, Prop. 4.3.6, Cor. 4.3.7, Prop. 4.3.8, Thm. 4.4.3]; see also Theorem 5 in Step 2. The finite projectivity of $H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},W_{A})$ over $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$ and its base-change property then follow from the fact that $\mathcal{R}_{A}^{\infty}(\Gamma)$ is a Fréchet–Stein algebra, finite flat over $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$, and that $H^{1}\bigl(C^{\bullet}_{\psi}(\mathscr{D}_{A})\bigr)$ is a coadmissible (hence finite projective) $\mathcal{R}_{A}^{\infty}(\Gamma)$–module; see [19, Lem. 4.3.4]. ∎
@@
(3.2) Classical Perrin–Riou big logarithm.
We recall the local Perrin–Riou regulator for a single de Rham Panchishkin representation.
Theorem 7 (Perrin–Riou big logarithm for a de Rham Panchishkin representation).
Let $V$ be a finite-dimensional $\mathbb{Q}_{p}$–vector space with a continuous de Rham $G_{\mathbb{Q}_{p}}$–action, and let $V^{+}\subset V$ be a Panchishkin subspace in the usual sense (cf. Panchishkin [3] or [16]). Then there exists a canonical $\Lambda(\Gamma)$–linear map
$\mathcal{L}_{V,V^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\otimes_{\mathbb{Q}_{p}}D_{\mathrm{cris}}(V)$
with the following interpolation property: for every integer twist $V(j)$ in the Bloch–Kato range determined by $V^{+}$ and every continuous character $\chi\colon\Gamma\to\overline{\mathbb{Q}}_{p}^{\times}$ of finite order, the specialisation of $\mathcal{L}_{V,V^{+}}$ at $\chi$ is, up to an explicit non-zero scalar depending only on $V$ and $\chi$, the Bloch–Kato dual exponential or logarithm map for $V(j)$.

### v2

@@ 2.2.3 Finiteness and base change for $(\varphi,\Gamma)$-cohomology
    The Iwasawa cohomology complex $C^{\bullet}_{\psi}(\mathsf{M})$ lies in $D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}_{A}^{\infty}(\Gamma)\bigr)$, and the $(\varphi,\Gamma)$-cohomology complex $C^{\bullet}_{\varphi,\Gamma}(\mathsf{M})$ lies in $D^{-}_{\mathrm{perf}}(A)$ [19, Thm. 4.4.1, Thm. 4.4.2].
@@
    $C^{\bullet}_{\psi}(\mathsf{M})\otimes^{\mathbf{L}}_{\mathcal{R}_{A}^{\infty}(\Gamma)}\mathcal{R}_{B}^{\infty}(\Gamma)\longrightarrow C^{\bullet}_{\psi}\bigl(\mathsf{M}\widehat{\otimes}_{A}B\bigr),$
@@
We apply Theorem 2.5 with $A=\mathcal{O}(U)$ and $\mathsf{M}=\mathscr{D}$ or $\mathsf{M}=\mathscr{D}^{+}$. This shows that
$C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}),\ C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}^{+})\in D^{-}_{\mathrm{perf}}\bigl(\mathcal{O}(U)\bigr),$
@@
$C^{\bullet}_{\psi}(\mathscr{D}),\ C^{\bullet}_{\psi}(\mathscr{D}^{+})\in D^{-}_{\mathrm{perf}}\bigl(\mathcal{R}_{\mathcal{O}(U)}^{\infty}(\Gamma)\bigr).$
Moreover, for any morphism of affinoid algebras $\mathcal{O}(U)\to B$ (for instance $B=\mathcal{O}(U^{\prime})$ for an affinoid subdomain $U^{\prime}\subset U$, or $B=k(x)$ for the residue field at a point $x\in U$), the base-change isomorphisms of Theorem 2.5(ii) give quasi-isomorphisms
$C^{\bullet}_{\varphi,\Gamma}(\mathscr{D})\otimes^{\mathbf{L}}_{\mathcal{O}(U)}B\xrightarrow{\ \sim\ }C^{\bullet}_{\varphi,\Gamma}\bigl(\mathscr{D}\widehat{\otimes}_{\mathcal{O}(U)}B\bigr),$
@@
$C^{\bullet}_{\psi}(\mathscr{D})\otimes^{\mathbf{L}}_{\mathcal{R}_{\mathcal{O}(U)}^{\infty}(\Gamma)}\mathcal{R}_{B}^{\infty}(\Gamma)\xrightarrow{\ \sim\ }C^{\bullet}_{\psi}\bigl(\mathscr{D}\widehat{\otimes}_{\mathcal{O}(U)}B\bigr),$
@@
Remark 2.6.
[…]
$\mathcal{L}_{V_{U},V_{U}^{+}}:H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Z}_{p}}\mathcal{O}(U),$
and to relate its specialisations at classical points to the Bloch–Kato exponentials.
2.3 Step 3: Construction of the big logarithm and the finite-slope regulator
In this final step we fix once and for all the affinoid neighbourhood $U$ and the triangulated $(\varphi,\Gamma)$-module
$(\mathscr{D},\mathscr{D}^{+})$
over $U$ constructed in Step 2. Thus $V_{U}$ is the rank four finite projective $\mathcal{O}(U)$-module equipped with a continuous $\mathcal{O}(U)$-linear action of $G_{\mathbb{Q}}$ whose specialisations are the Rankin–Selberg Galois representations $V_{x}$, and
$\mathscr{D}=D_{\operatorname{rig}}^{\dagger}\bigl(V_{U}\big|_{G_{\mathbb{Q}_{p}}}\bigr)$
[…]
We first introduce Iwasawa-theoretic notation and recall the local Perrin–Riou big logarithm for a single Panchishkin representation. We then explain how to extend it to the family $(V_{U},V_{U}^{+})$ using relative $(\varphi,\Gamma)$-modules and the Perrin–Riou formalism, and finally we state the Euler–system conjecture and the resulting conditional construction of $L_{p}^{\mathrm{fs}}$.
2.3.1 Iwasawa cohomology and the distribution algebra
@@
denote the algebra of $\mathbb{Q}_{p}$-valued locally analytic distributions on $\Gamma$, i.e. the completion of $\Lambda_{\mathbb{Q}_{p}}(\Gamma)$ in its natural Fréchet topology (cf. [30, §2.2] and [31, §2.1]). Equivalently, $\mathcal{H}(\Gamma)$ is the strong continuous dual of the space of locally analytic $\mathbb{Q}_{p}$-valued functions on $\Gamma$.
For any $p$-adic representation $W$ of $G_{\mathbb{Q}_{p}}$ on a finite-dimensional $\mathbb{Q}_{p}$-vector space, its (cyclotomic) Iwasawa cohomology is defined by
$H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W):=H^{i}\bigl(\mathbb{Q}_{p},W\otimes_{\mathbb{Q}_{p}}\Lambda(\Gamma)^{\vee}\bigr)\qquad(i\geq 0),$
where $\Lambda(\Gamma)^{\vee}:=\operatorname{Hom}_{\mathrm{cts}}(\Lambda(\Gamma),\mathbb{Q}_{p})$ with the natural $G_{\mathbb{Q}_{p}}$-action; see [18, App. A.2–A.4] and Greenberg [28]. By construction, $\Lambda(\Gamma)$ acts on $\Lambda(\Gamma)^{\vee}$ via the right regular representation, and hence $H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)$ carries a natural $\Lambda(\Gamma)$-module structure for every $i$.
Lemma 2.7 (Finiteness of local Iwasawa cohomology).
If $W$ is a finite-dimensional $\mathbb{Q}_{p}$-representation of $G_{\mathbb{Q}_{p}}$, then $H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)$ is a finitely generated $\Lambda(\Gamma)$-module for all $i\geq 0$. In particular, $H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)$ is finitely generated over $\Lambda(\Gamma)$.
@@
Choose a $G_{\mathbb{Q}_{p}}$-stable $\mathbb{Z}_{p}$-lattice $T\subset W$, and write $K_{\infty}=\mathbb{Q}_{p}(\mu_{p^{\infty}})$. Following [18, App. A.2], let
@@
with respect to the corestriction maps; then $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ is naturally a $\Lambda(\Gamma)$-module, and there is a canonical identification
$Z^{i}_{\infty}(\mathbb{Q}_{p},T)\cong H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},T):=H^{i}\bigl(\mathbb{Q}_{p},T\otimes_{\mathbb{Z}_{p}}\Lambda(\Gamma)^{\vee}\bigr)$
for all $i$ (see [18, App. A.2–A.3]). Moreover, Proposition A.2.3 of loc. cit. shows that $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ is a finitely generated $\Lambda(\Gamma)$-module for $i=0,1,2$, and $Z^{i}_{\infty}(\mathbb{Q}_{p},T)=0$ for $i\geq 3$. Since $W=T\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$, we have
$H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)\cong H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},T)\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}\cong Z^{i}_{\infty}(\mathbb{Q}_{p},T)\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$
as $\Lambda(\Gamma)$-modules, and the finite generation of $Z^{i}_{\infty}(\mathbb{Q}_{p},T)$ implies the finite generation of $H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)$ for all $i$. ∎
In order to apply Perrin–Riou’s regulator, we pass from $\Lambda(\Gamma)$-coefficients to the algebra $\mathcal{H}(\Gamma)$ of locally analytic distributions by scalar extension along the canonical map $\Lambda(\Gamma)\to\mathcal{H}(\Gamma)$ and set
$H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)_{\mathcal{H}}:=H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)\widehat{\otimes}_{\Lambda(\Gamma)}\mathcal{H}(\Gamma).$
When no confusion can arise, we continue to denote this $\mathcal{H}(\Gamma)$-module simply by $H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W)$.
If $A$ is a $\mathbb{Q}_{p}$-affinoid algebra and $W_{A}$ is a finite projective $A$-module with a continuous $A$-linear $G_{\mathbb{Q}_{p}}$-action, we define the family Iwasawa cohomology by
$H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{A}):=H^{i}\bigl(\mathbb{Q}_{p},W_{A}\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)^{\vee}\bigr)\qquad(i\geq 0),$
which is a priori an $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$-module. Let $\mathcal{R}_{A}^{\infty}(\Gamma)$ denote the Fréchet–Stein Iwasawa algebra of $\Gamma$ over $A$ considered in [19, §4.2]; there is a natural finite flat homomorphism
@@
The relative $(\varphi,\Gamma)$-module formalism of Kedlaya–Pottharst–Xiao gives a more precise description.
Proposition 2.8 (Relative local Iwasawa cohomology via $(\varphi,\Gamma)$-modules).
Let $A$ be a reduced $\mathbb{Q}_{p}$-affinoid algebra and $W_{A}$ a finite projective $A$-module with a continuous $A$-linear $G_{\mathbb{Q}_{p}}$-action. Let $\mathscr{D}_{A}:=D_{\operatorname{rig}}^{\dagger}(W_{A})$ be the associated family of $(\varphi,\Gamma)$-modules over the relative Robba ring $\mathcal{R}_{A}$. Then:
@@
    $H^{i}\bigl(C^{\bullet}_{\psi}(\mathscr{D}_{A})\bigr)\cong H^{i}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{A})\widehat{\otimes}_{\Lambda(\Gamma)}\mathcal{R}_{A}^{\infty}(\Gamma)$
@@
    in particular, $H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{A})$ is a finite projective $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$-module, and for every morphism of affinoids $A\to B$ the canonical map
    $H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{A})\widehat{\otimes}_{A\widehat{\otimes}\Lambda(\Gamma)}B\widehat{\otimes}\Lambda(\Gamma)\xrightarrow{\ \sim\ }H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{B})$
@@
The existence, perfectness, and base-change properties of the complexes $C^{\bullet}_{\varphi,\Gamma}(\mathscr{D}_{A})$ and $C^{\bullet}_{\psi}(\mathscr{D}_{A})$, as well as the identifications of their cohomology with Galois and Iwasawa cohomology after extension of scalars to $\mathcal{R}_{A}^{\infty}(\Gamma)$, are proved in [19, Rem. 4.3.3, Prop. 4.3.6, Cor. 4.3.7, Prop. 4.3.8, Thm. 4.4.3]. The finite projectivity of $H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},W_{A})$ over $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$ and its base-change property then follow from the fact that $\mathcal{R}_{A}^{\infty}(\Gamma)$ is a Fréchet–Stein algebra, finite flat over $A\widehat{\otimes}_{\mathbb{Q}_{p}}\Lambda(\Gamma)$, and that $H^{1}\bigl(C^{\bullet}_{\psi}(\mathscr{D}_{A})\bigr)$ is a coadmissible (hence finite projective) $\mathcal{R}_{A}^{\infty}(\Gamma)$-module; see [19, Lem. 4.3.4]. ∎
@@
2.3.2 Classical Perrin–Riou big logarithm
We briefly recall the local Perrin–Riou regulator for a single de Rham Panchishkin representation.
Theorem 2.9 (Perrin–Riou big logarithm).
Let $V$ be a finite-dimensional $\mathbb{Q}_{p}$-vector space with a continuous de Rham $G_{\mathbb{Q}_{p}}$-action, and let $V^{+}\subset V$ be a Panchishkin subspace in the usual sense (cf. Panchishkin [3] or [16]). Then there exists a canonical $\Lambda(\Gamma)$-linear map
$\mathcal{L}_{V,V^{+}}:H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},V^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\otimes_{\mathbb{Q}_{p}}D_{\mathrm{cris}}(V)$
with the following interpolation property: for every integer twist $V(j)$ in the Bloch–Kato range determined by $V^{+}$ and every continuous character $\chi:\Gamma\to\overline{\mathbb{Q}}_{p}^{\times}$ of finite order, the specialisation of $\mathcal{L}_{V,V^{+}}$ at $\chi$ is, up to an explicit non-zero scalar depending only on $V$ and $\chi$, the Bloch–Kato dual exponential or logarithm map for $V(j)$.